# Entropic Trade-off Relations in Stochastic Thermodynamics via Replica Markov Processes

### v2

@@ II.4 Entropic bound on trajectory observable
We consider the uncertainty in trajectory observables. Trajectory observables play a fundamental role in uncertainty relations in stochastic thermodynamics (Fig. 2(a)). Consider a general observable $N_{\bullet}(\Gamma)$. Using Eq. (14), we obtain the upper bound on the time derivative of the Tsallis entropy (see Appendix A for the derivation):
$\left|\frac{\partial}{\partial\tau}H_{q}^{\mathrm{T}}\left[P\left(N_{\bullet}\right)\right]\right|\leq\frac{\sqrt{q\mathcal{A}(\tau)}}{2\tau(q-1)},$ (24)
where $q=2,3,\cdots$. The bound in Eq. (24) provides the upper bound of the change in the Tsallis entropy. Equation (24) shows that the change in entropy is larger for larger $\mathcal{A}(\tau)$. By performing the integration in Eq. (24), we can derive the upper bound on the Tsallis entropy with respect to the observable $N_{\circ}$ (see Appendix A for the derivation):
@@
Outside the range of Eq. (26), the bound of Eq. (25) provides a trivial upper bound 1/(q1)1/(q-1), which directly follows from Eq. (21). On the right-hand side of Eq. (25), the integral $\int_{0}^{\tau}dt\,\sqrt{\mathcal{A}(t)}/t$ often appears in thermodynamic trade-off relations [44]. We have demonstrated the consequence of considering Eq. (14) in the replica scenario. We can derive another trade-off relation from Eq. (17). From Eq. (17), we obtain another trade-off relation (see Appendix C for the derivation):
@@
Both left-hand sides of Eqs. (25) and (28) are Tsallis entropies, and the right-hand sides involve the dynamical activity. In a general transient setting, it is unclear which of the two is tighter; however, in the steady state, Eq. (28) provides the tighter bound.
Traditionally, uncertainty in trajectory observables has been characterized using variance $\mathrm{Var}[N]$, whose relative value is known to be bounded from below by the inverse of the dynamical activity. For a steady-state Markov process and the counting observable $N(\Gamma)$ [Eq. (5)], it is known that the following relation holds [3, 5]:
$\frac{\mathrm{Var}[N]}{\mathbb{E}[N]^{2}}\geq\frac{1}{\mathfrak{a}_{\mathrm{ss}}\tau},$ (29)
where $\mathfrak{a}_{\mathrm{ss}}$ is the steady-state dynamical activity. Moreover, for non-steady-state conditions and for the observable $N_{\circ}(\Gamma)$ satisfying Eq. (13), the following relation is known to hold [45, 49]:
@@
which follows directly from Eq. (17) with $K=1$. Our results show that this fundamental constraint extends beyond relative variance: the uncertainty measure based on Tsallis (or Rényi) entropy is also limited by the dynamical activity. However, there is an important difference between Eq. (28) [or Eq. (25)] and Eq. (29). When the variance $\mathrm{Var}[N]$ decreases, the uncertainty in $N$ becomes smaller; thus Eq. (29) provides a lower bound on the uncertainty in $N$. By contrast, Eq. (28) [or Eq. (25)] yields an upper bound on the uncertainty in the trajectory observable. Therefore, these two relations give complementary lower and upper bounds on the uncertainty of $N$, each expressed in terms of different quantities. The Tsallis entropy reduces to the Shannon entropy for $q\to 1$; however, the bound in Eq. (28) is not defined for $q=1$. Regarding the Shannon entropy for trajectory observables, Ref. [50] derived a lower bound for the Shannon entropy using the entropy production. The bound in Eq. (28) provides the upper bound of the Tsallis entropy and the bound comprises the dynamical activity.
@@ II.5 Entropic bound on state distribution
So far we are concerned with entropic trade-off relations for the trajectory observables $N_{\bullet}(\Gamma)$ and $N_{\circ}(\Gamma)$. Next, we derive entropic trade-off relations for the state probability distribution $p_{\mu}(t)$. The entropy of the state distribution can quantify the extent of diffusion in a network. Specifically, let us consider a random walker on a network whose dynamics is governed by a continuous time Markov process [Eq. (3)], which is extensively investigated in the literature [51, 52, 53]. It has many applications in web search engines [54], community detection [55], embedding algorithms [56], and computational biology [57], to name but a few. Here, our objective is to quantify the extent of diffusion in the network (Fig. 2(b)). Typically, the mean squared displacement, mixing time, or the spectral gap of the generator are used [51, 53]. Entropy is another metric that is used to quantify diffusion in a network. The diversity entropy [58, 59] is defined as the entropy of the probability of position of the walker and measures how widely a random walker can spread over the network. The entropy rate [60] quantifies how quickly the uncertainty increases as the random walk evolves.
The degree of diffusion can be quantified by the Tsallis entropy (or Rényi) $H_{q}^{\mathrm{T}}[p_{\mu}(\tau)]$ (or $H_{\alpha}^{\mathrm{R}}[p_{\mu}(\tau)]$). When this entropy is zero, it indicates that no propagation has occurred in the network, and as the degree of propagation increases, the value of this entropy increases. From Eq. (24), we obtain
@@
where $q=2,3,\cdots$. The derivation of Eq. (31) is shown in Appendix A. As expected, a larger dynamical activity allows for faster diffusion, which is quantified by the change in the Tsallis entropy. By integrating Eq. (31), we obtain the upper bound on the Tsallis entropy of the state distribution:
$\left|H_{q}^{\mathrm{T}}[p_{\mu}(\tau)]-H_{q}^{\mathrm{T}}[p_{\mu}(0)]\right|\leq\frac{\sqrt{q}}{2(q-1)}\int_{0}^{\tau}\frac{\sqrt{\mathcal{A}(t)}}{t}dt,$ (32)
where $q=2,3,\cdots$. Equation (32) describes the difference of the Tsallis entropy at $t=0$ and $t=\tau$. The bounds in Eqs. (31) and (32) hold for arbitrary initial state $p_{\mu}(0)$. When the initial state is $B_{\mu_{0}}$ (i.e., $p_{\mu_{0}}(0)=1$), then the tighter relation holds (see Appendix A for the derivation):
@@
where $q=2,3,\cdots$ and the condition given in Eq. (26) should be satisfied. As in the trajectory observable case, we can derive another trade-off relation from Eq. (17). Using Eq. (17), the following relation holds for the Rényi entropy (see Appendix C):
@@
where $q=2,3,\cdots$. The left-hand sides of Eqs. (33) and (35) are the same, and in both equations the upper bound of the Tsallis entropy at time $\tau$ is given by the dynamical activity; a larger dynamical activity allows for a larger Tsallis entropy. On the other hand, while Eq. (33) requires information beyond the neighborhood of the initial node, Eq. (35) does not. Specifically, in Eq. (35), it is bounded from above by $\tau$ and the local escape rate from the initial node $B_{\mu_{0}}$, which is the sum of the transition rates from the initial node $B_{\mu_{0}}$ to the nodes that can be reached within a jump ( note that $\sum_{\mu(\mu\neq\mu_{0})}W_{\mu\mu_{0}}=\mathfrak{a}(t=0)$ corresponds to the dynamical activity at the initial time). Surprisingly, this right-hand side of Eq. (35) depends only on the time $\tau$ and on the local escape rate from $B_{\mu_{0}}$ (see Fig. 3). In other words, regardless of how large the overall network is, we can compute the upper bound of the entropy using only information about the initial node and the transition rates to nodes that are reachable in a single hop.
@@ II.6 Bound on extreme value
By using replicas, we can consider more general observables. As the next example, we derive a trade-off relation for the extreme-value of a stochastic process. Extreme-value statistics studies the behavior of the largest (and similarly, the smallest) value in a collection of random variables $\{X_{k}\}$; that is, it focuses on the random variable $\max_{k}X_{k}$ (and analogously $\min_{k}X_{k}$):
$X^{M\max}\equiv\max_{1\leq k\leq M}X_{k}.$ (36)
Consider the observable $N_{\circ}^{M,\max}$, which is the maximum of $N_{\circ}^{(k)}(\Gamma^{(k)})$:
@@
When the state of the Markov process describes the position, $N^{M\max}_{\circ}$ expresses the maximum position at time $t=\tau$. In reliability engineering, the system load can be modeled using a Markov process [61, 62]. In this case, the maximum value of the stochastic process represents the maximum load of the parallel system. Since $N_{\circ}^{M\max}$ satisfies Eq. (6), using Eq. (18), we obtain
@@
Since the dynamical activity is multiplied by $M$, the relative variance of the maximum observable $N_{\circ}^{M\max}$ can be smaller than those of a single process. Usually, when we take the max operation, the mean becomes larger and the variance becomes smaller than those of the original distribution. For example, consider two independent random variables, each following a standard normal distribution. If we take the maximum of these two variables, the resulting variable will have a mean of $1/\sqrt{\pi}$ and a variance of $1-1/\pi$. Its mean is larger, and its variance is smaller when compared to the original standard normal distribution.
@@
$\left|\frac{\partial}{\partial\tau}H_{q}^{\mathrm{T}}[P(N_{\circ}^{M\max})]\right|\leq\frac{\sqrt{qM\mathcal{A}(\tau)}}{2\tau(q-1)},$ (39)
@@ II.7 Quantum replica Markov process
where the subscript Q``Q" is used to emphasize that $\Gamma_{Q}$ is a quantum trajectory. In what follows, variables and functions with subscript $Q$ denote those in the quantum system. The sequence $\Gamma_{Q}$ characterizes one realization of the monitored dynamics and is the basic object from which we construct trajectory-dependent observables. Consider a trajectory observable $N_{Q}(\Gamma_{Q})$, which is a quantum generalization of Eq. (5):
$N_{Q}(\Gamma_{Q})=\sum_{m}C_{m}N_{m},$ (46)
where $N_{m}$ is the number of $m$-th jumps within $0,\tau]$ and $C_{m}$ denotes the corresponding real weight. Let $\Gamma_{Q\varnothing}$ be a quantum trajectory in which there is no jump within the interval $0,\tau]$. Analogously to the condition of Eq. (13), we define the observable $N_{Q\circ}(\Gamma_{Q})$ that satisfies the following condition:
$N_{Q\circ}(\Gamma_{Q\varnothing})=0.$ (47)
$N_{{Q}\circ}$ includes counting observables, which simply count the number of jumps. Let $N_{Q\bullet}(\Gamma_{Q})$ be an arbitrary function defined on $\Gamma_{Q}$. If we denote the sets of functions $N_{Q}$, $N_{Q\circ}$, and $N_{Q\bullet}$ by $\{N_{Q}\}$, $\{N_{Q\circ}\}$, and $\{N_{Q\bullet}\}$, respectively, then inclusion holds:
$\{N_{Q}\}\subset\{N_{Q\circ}\}\subset\{N_{Q\bullet}\}.$ (48)
In a similar manner to the classical case, we consider $K$ replicas of the identical quantum Markov process. Specifically, $\Gamma_{Q}^{(\otimes K)}$ can be constructed from $\left(\Gamma_{Q}^{(1)},\Gamma_{Q}^{(2)},\cdots,\Gamma_{Q}^{(K)}\right)$, where $\Gamma_{Q}^{(k)}$ is a trajectory in the $k$-th process.
@@ II.8 Quantum replica trade-off relations
As in the classical cases, it is straightforward to derive trade-off relations in the quantum replica process. Specifically, it is known that the following relation holds [44]:
$\frac{\mathrm{Var}[N_{Q\mathrm{\bullet}}]}{\tau^{2}\left(\partial_{\tau}\mathbb{E}[N_{Q\bullet}]\right)^{2}}\geq\frac{1}{\mathcal{B}(\tau)}.$ (49)
@@
$\frac{\mathrm{Var}[N_{Q}]}{\mathbb{E}[N_{Q}]^{2}}\geq\frac{1}{\mathcal{B}(\tau)}$ (50)
which holds for the counting observable $N_{Q}(\Gamma_{Q})$ defined in Eq. (46). In Eqs. (49) and (50), $\mathcal{B}(\tau)$ denotes the quantum dynamical activity within the interval $0,\tau] [$17, 44, 64]. The classical dynamical activity [Eq. (16)] quantifies the extent of activity of the dynamics. For classical Markov processes, the state change occurs only when there are jumps. However, for the quantum scenario, the state change can occur even without jumps due to the coherent contribution arising from $i[H,\rho]$ term in Eq. (43). The quantum dynamical activity plays a central role in trade-off relations in quantum thermodynamics [17, 44, 65, 64, 66]. For the explicit expression of $\mathcal{B}(\tau)$, see Appendix F.
@@
$\frac{\mathrm{Var}[N_{Q\mathrm{\bullet}}^{(\otimes K)}]}{\tau^{2}\left(\partial_{\tau}\mathbb{E}[N_{Q\bullet}^{(\otimes K)}]\right)^{2}}\geq\frac{1}{K\mathcal{B}(\tau)}.$ (51)
Similarly to the classical counterpart [Eq. (14)], the right-hand side of Eq. (51) depends on the quantum dynamical activity $\mathcal{B}(\tau)$. The left-hand side of Eq. (51) comprises the observable $N_{Q\bullet}(\Gamma_{Q})$, which is an arbitrary function of a quantum trajectory $\Gamma_{Q}$. Moreover, for the observable $N_{Q\circ}$, which requires the condition given in Eq. (47), the following relation holds (see Appendix D for details):
@@
The derivation of Eq. (52) is based on the bound proved in Ref. [45]. For the classical case, the right-hand side depends solely on the initial dynamical activity $\mathfrak{a}(t=0)$. On the other hand, $\mathcal{B}(\tau)$ depends on the dynamics within the interval $0,\tau]$.

### v3

@@ II.4 Entropic bound on trajectory observable
We consider the uncertainty in trajectory observables. Trajectory observables play a fundamental role in uncertainty relations in stochastic thermodynamics (Fig. 2(a)). Consider a general observable $N(\Gamma)$. Using Eq. (14), we obtain the upper bound on the time derivative of the Tsallis entropy (see Appendix A for the derivation):
$\left|\frac{\partial}{\partial\tau}H_{q}^{\mathrm{T}}\left[P\left(N\right)\right]\right|\leq\frac{\sqrt{q\mathcal{A}(\tau)}}{2\tau(q-1)},$ (24)
where $q=2,3,\cdots$. The bound in Eq. (24) provides the upper bound on the change in the Tsallis entropy. Equation (24) shows that the change in entropy is larger for larger $\mathcal{A}(\tau)$. Integrating in Eq. (24), we can derive the upper bound on the Tsallis entropy with respect to the observable $N_{\circ}$ (see Appendix A for the derivation):
@@
Outside the range of Eq. (26), the bound of Eq. (25) provides a trivial upper bound 1/(q1)1/(q-1), which directly follows from Eq. (20). On the right-hand side of Eq. (25), the integral $\int_{0}^{\tau}dt\,\sqrt{\mathcal{A}(t)}/t$ often appears in thermodynamic trade-off relations [46].
We have demonstrated the consequence of considering Eq. (14) in the replica scenario. We can derive another trade-off relation from Eq. (17). From Eq. (17), we obtain another trade-off relation (see Appendix D for the derivation):
@@
Both left-hand sides of Eqs. (25) and (28) are Tsallis entropies, and the right-hand sides involve the dynamical activity. In a general transient setting, it is unclear which of the two is tighter; however, in the steady state, Eq. (28) provides a tighter bound.
Traditionally, uncertainty in trajectory observables has been characterized using the variance $\mathrm{Var}[N_{\sharp}]$, whose relative value is known to be bounded from below by the inverse of the dynamical activity. For a steady-state Markov process and the counting observable $N_{\sharp}(\Gamma)$ [Eq. (6)], it is known that the following relation holds [3, 5]:
$\frac{\mathrm{Var}[N_{\sharp}]}{\mathbb{E}[N_{\sharp}]^{2}}\geq\frac{1}{\mathfrak{a}_{\mathrm{ss}}\tau},$ (29)
where $\mathfrak{a}_{\mathrm{ss}}$ is the steady-state dynamical activity. Moreover, under non-steady-state conditions and for the observable $N_{\circ}(\Gamma)$ satisfying Eq. (13), the following relation is known to hold [47, 51]:
@@
which follows directly from Eq. (17) with $K=1$. Our results show that this fundamental constraint extends beyond relative variance: the uncertainty measure based on Tsallis (or Rényi) entropy is also limited by the dynamical activity. However, there is an important difference between Eq. (28) [or Eq. (25)] and Eq. (29). When the variance $\mathrm{Var}[N_{\sharp}]$ decreases, the uncertainty in $N_{\sharp}$ becomes smaller; thus, Eq. (29) provides a lower bound on the uncertainty in $N_{\sharp}$. By contrast, Eq. (28) [or Eq. (25)] yields an upper bound on the uncertainty in the trajectory observable. Therefore, these two relations provide complementary lower and upper bounds on the uncertainty of the trajectory observable, each expressed in terms of different quantities. The Tsallis entropy reduces to the Shannon entropy for $q\to 1$; however, the bound in Eq. (28) is not defined for $q=1$. Regarding the Shannon entropy for trajectory observables, Ref. [52] derived a lower bound for the Shannon entropy using the entropy production. The bound in Eq. (28) provides an upper bound on the Tsallis entropy, and this bound involves the dynamical activity.
@@ II.5 Entropic bound on state distribution
So far, we have been concerned with entropic trade-off relations for the trajectory observables. Next, we derive entropic trade-off relations for the state probability distribution $p_{\mu}(t)$. The entropy of the state distribution can quantify the extent of diffusion in a network. Suppose that a random walker moves on a network whose dynamics is governed by a continuous-time Markov process [Eq. (3)], which has been extensively investigated in the literature [53, 54, 55]. It has many applications in web search engines [56], community detection [57], embedding algorithms [58], and computational biology [59], to name but a few. Here, our objective is to quantify the extent of diffusion in the network (Fig. 2(b)). Typically, the mean squared displacement, mixing time, or the spectral gap of the generator is used [53, 55]. Entropy is another metric used to quantify diffusion in a network. The diversity entropy [60, 61] is defined as the entropy of the probability distribution of the position of the walker and measures how widely a random walker can spread over the network. The entropy rate [62] quantifies how quickly the uncertainty increases as the random walk evolves.
The degree of diffusion can be quantified by the Tsallis entropy (or Rényi entropy) $H_{q}^{\mathrm{T}}[p_{\mu}(\tau)]$ (or $H_{\alpha}^{\mathrm{R}}[p_{\mu}(\tau)]$). When this entropy is zero, it indicates that no propagation has occurred in the network, and as the degree of propagation increases, the value of this entropy increases. From Eq. (24), we obtain
@@
where $q=2,3,\cdots$. The derivation of Eq. (31) is shown in Appendix B. As expected, larger dynamical activity allows for faster diffusion, which is quantified by the change in the Tsallis entropy. By integrating Eq. (31), we obtain an upper bound on the difference in the Tsallis entropy of the state distribution:
$\left|\Delta H_{q}^{\mathrm{T}}[p_{\mu}]\right|\leq\frac{\sqrt{q}}{2(q-1)}\int_{0}^{\tau}\frac{\sqrt{\mathcal{A}(t)}}{t}dt,$ (32)
where $q=2,3,\cdots$ and $\Delta H_{q}^{\mathrm{T}}[p_{\mu}]\equiv H_{q}^{\mathrm{T}}[p_{\mu}(\tau)]-H_{q}^{\mathrm{T}}[p_{\mu}(0)]$. A similar relation was obtained for $q=2$ in Ref. [63]. Equation (32) describes the difference in the Tsallis entropy at $t=0$ and $t=\tau$. In the Landauer principle [64, 65], the following relation holds:
$\Delta H_{1}^{\mathrm{T}}[p_{\mu}]\geq-\beta\Delta Q(\tau),$ (33)
where $\beta$ is the inverse temperature and $H_{1}^{\mathrm{T}}[p_{\mu}]\equiv\lim_{q\rightarrow 1}H_{q}^{\mathrm{T}}[p_{\mu}]=-\sum_{\mu}p_{\mu}\ln p_{\mu}$ corresponds to the Shannon entropy. In Eq. (33), the change in the Shannon entropy $H_{1}^{\mathrm{T}}\left[p_{\mu}(t)\right]$ from the initial distribution $p_{\mu}(0)$ to the distribution at time $\tau$, $p_{\mu}(\tau)$, is constrained by the heat exchanged with a thermal reservoir. With the sign convention that $\Delta Q(\tau)$ is the heat absorbed by the reservoir up to time $\tau$, this inequality relates the information change to the thermodynamic cost. In contrast, Eq. (32) provides an upper bound on the change in Tsallis entropy, where the upper bound comprises the dynamical activity.
The bounds in Eqs. (31) and (32) hold for an arbitrary initial state $p_{\mu}(0)$. When the initial state is $B_{\mu_{0}}$ (i.e., $p_{\mu_{0}}(0)=1$), the tighter relation holds (see Appendix B for the derivation):
@@
where $q=2,3,\cdots$, and the condition given in Eq. (26) should be satisfied.
As in the trajectory observable case, we can derive another trade-off relation from Eq. (17). Using Eq. (17), the following relation holds for the Rényi entropy (see Appendix D):
@@
where $q=2,3,\cdots$. The left-hand sides of Eqs. (34) and (36) are the same, and, in both equations, the upper bound of the Tsallis entropy at time $\tau$ is given by the dynamical activity; a larger dynamical activity allows for a larger Tsallis entropy. On the other hand, while Eq. (34) requires information beyond the neighborhood of the initial node, Eq. (36) does not. Specifically, in Eq. (36), it is bounded from above by $\tau$ and the local escape rate from the initial node $B_{\mu_{0}}$, which is the sum of the transition rates from the initial node $B_{\mu_{0}}$ to the nodes that can be reached in a single jump (note that $\sum_{\mu(\mu\neq\mu_{0})}W_{\mu\mu_{0}}=\mathfrak{a}(t=0)$ corresponds to the dynamical activity at the initial time). Therefore, the right-hand side of Eq. (36) depends only on the time $\tau$ and the local escape rate from $B_{\mu_{0}}$ (see Fig. 3). In other words, regardless of how large the overall network is, we can compute the upper bound of the entropy using only information about the initial node and the transition rates to nodes that are reachable in a single hop.
@@ II.6 Bound on extreme value
By using replicas, we can consider more general observables. As the next example, we derive a trade-off relation for the extreme value of a stochastic process. Extreme-value statistics studies the behavior of the largest (and similarly, the smallest) value in a collection of random variables $\{X_{k}\}$; that is, it focuses on the random variable $\max_{k}X_{k}$ (and analogously $\min_{k}X_{k}$). Consider the observables $N^{M,\max}$ and $N_{\circ}^{M,\max}$, which are the maximum of $N^{(k)}(\Gamma^{(k)})$ and $N_{\circ}^{(k)}(\Gamma^{(k)})$, respectively:
$N^{M\max}(\Gamma^{(\otimes M)})$ $\equiv\max_{1\leq k\leq M}N^{(k)}(\Gamma^{(k)}),$ (37)
@@
When the state of the Markov process describes the position, $N^{M\max}$ and $N^{M\max}_{\circ}$ express the maximum position at time $t=\tau$ in $M$ processes. In reliability engineering, the system load can be modeled using a Markov process [66, 67]. In this case, the maximum value of the stochastic process represents the maximum load of the parallel system. Since $N^{(\otimes K)}$ in Eq. (14) allows for an arbitrary observable, the following relation holds:
$\frac{\mathrm{Var}[N^{M\max}]}{\tau^{2}\left(\partial_{\tau}\mathbb{E}[N^{M\max}]\right)^{2}}\geq\frac{1}{M\mathcal{A}(\tau)}.$ (39)
Moreover, since $N_{\circ}^{M\max}$ satisfies Eq. (5), using Eq. (18), we obtain
@@
Since the dynamical activity is multiplied by $M$, the relative variance of the maximum observable $N_{\circ}^{M\max}$ can be smaller than that of a single process. Usually, when we take the max operation, the mean becomes larger and the variance becomes smaller than those of the original distribution. For example, consider two independent random variables, each following a standard normal distribution. If we take the maximum of these two variables, the resulting variable will have a mean of $1/\sqrt{\pi}$ and a variance of $1-1/\pi$. Its mean is larger, and its variance is smaller, when compared to the original standard normal distribution.
@@
$\left|\frac{\partial}{\partial\tau}H_{q}^{\mathrm{T}}[P(N^{M\max})]\right|\leq\frac{\sqrt{qM\mathcal{A}(\tau)}}{2\tau(q-1)},$ (41)
@@ II.7 Quantum replica Markov process
where the subscript Q``Q" is used to emphasize that $\Gamma_{Q}$ is a quantum trajectory. In what follows, variables and functions with the subscript $Q$ denote those in the quantum system. The sequence $\Gamma_{Q}$ characterizes one realization of the monitored dynamics and is the basic object from which we construct trajectory-dependent observables. We first introduce the most general class of trajectory observables. Let $N_{Q}(\Gamma_{Q})$ be an arbitrary real-valued function defined on $\Gamma_{Q}$. We denote the set of all such functions by $\{N_{Q}\}$. Next, we impose a natural constraint related to the no-jump trajectory. Let $\Gamma_{Q\varnothing}$ be a quantum trajectory in which there is no jump within $0,\tau]$. Analogous to the condition of Eq. (13), we define $N_{Q\circ}(\Gamma_{Q})$ as an observable satisfying
$N_{Q\circ}(\Gamma_{Q\varnothing})=0,$ (48)
and denote the set of such observables by $\{N_{Q\circ}\}$. Finally, we specify a particularly important subclass that depends only on jump counts with fixed weights. We define the counting-type trajectory observable $N_{Q\sharp}(\Gamma_{Q})$ by
$N_{Q\sharp}(\Gamma_{Q})\equiv\sum_{m}C_{m}N_{m},$ (49)
where $N_{m}$ is the number of $m$-th jumps within $0,\tau]$ and $C_{m}$ denotes the corresponding real weight. We denote the set of such observables by $\{N_{Q\sharp}\}$. By construction, $N_{Q\sharp}$ is included in $N_{Q\circ}$ (in particular, it vanishes on $\Gamma_{Q\varnothing}$). With these definitions, the inclusion relation is
$\{N_{Q\sharp}\}\subset\{N_{Q\circ}\}\subset\{N_{Q}\}.$ (50)
Analogous to the classical case, we consider $K$ identical replicas of the quantum Markov process. Specifically, $\Gamma_{Q}^{(\otimes K)}$ is constructed from the set $\left(\Gamma_{Q}^{(1)},\Gamma_{Q}^{(2)},\dots,\Gamma_{Q}^{(K)}\right)$, where each $\Gamma_{Q}^{(k)}$ represents a trajectory in the $k$-th replica.
@@ II.8 Quantum replica trade-off relations
As in the classical case, it is straightforward to derive trade-off relations in the quantum replica process. Specifically, it is known that the following relation holds [46]:
$\frac{\mathrm{Var}[N_{Q}]}{\tau^{2}\left(\partial_{\tau}\mathbb{E}[N_{Q}]\right)^{2}}\geq\frac{1}{\mathcal{B}(\tau)}.$ (51)
@@
$\frac{\mathrm{Var}[N_{Q\sharp}]}{\mathbb{E}[N_{Q\sharp}]^{2}}\geq\frac{1}{\mathcal{B}(\tau)}$ (52)
which holds for the counting observable $N_{Q\sharp}(\Gamma_{Q})$ defined in Eq. (49). In Eqs. (51) and (52), $\mathcal{B}(\tau)$ denotes the quantum dynamical activity within the interval $0,\tau] [$19, 46, 69]. The classical dynamical activity [Eq. (16)] quantifies the extent of activity in the dynamics. For classical Markov processes, the state changes occur only when there are jumps. However, in the quantum scenario, state changes can occur even without jumps due to the coherent contribution arising from the $i[H,\rho]$ term in Eq. (45). The quantum dynamical activity plays a central role in trade-off relations in quantum thermodynamics [19, 46, 70, 69, 71]. For the explicit expression of $\mathcal{B}(\tau)$, see Appendix G.
@@
$\frac{\mathrm{Var}[N_{Q}^{(\otimes K)}]}{\tau^{2}\left(\partial_{\tau}\mathbb{E}[N_{Q}^{(\otimes K)}]\right)^{2}}\geq\frac{1}{K\mathcal{B}(\tau)}.$ (53)
Similarly to the classical counterpart [Eq. (14)], the right-hand side of Eq. (53) depends on the quantum dynamical activity $\mathcal{B}(\tau)$. The left-hand side of Eq. (53) comprises the observable $N_{Q}^{(\otimes K)}(\Gamma_{Q}^{(\otimes K)}$, which is an arbitrary function of a quantum trajectory $\Gamma_{Q}^{(\otimes K)}$. Moreover, for the observable $N_{Q\circ}^{(\otimes K)}$, which requires the condition given in Eq. (48), the following relation holds using the result of Ref. [47] (see Appendix E for details):
@@
For the classical case, the right-hand side depends solely on the initial dynamical activity $\mathfrak{a}(t=0)$. On the other hand, $\mathcal{B}(\tau)$ depends on the dynamics within the interval $0,\tau]$.